OpenStudy (gabylovesyou):

The volume of a cylinder is 24 ft3. The volume of a cube is 54 ft3. What is the difference between the volume of a cone that fits exactly inside the cylinder and the volume of a pyramid that fits exactly inside the cube? Provide an explanation for your answer to receive full credit.

OpenStudy (gabylovesyou):

@jim_thompson5910

OpenStudy (gabylovesyou):

@paki

OpenStudy (paki):

do you know the formulas of volume of cube and cylinder...?

OpenStudy (gabylovesyou):

v = a^3 is volume of a cube ..

OpenStudy (gabylovesyou):

Volume of a cylinder:V = πr2h

OpenStudy (paki):

hmmm... use this relation here... Volume of a pyramid = 1/3 volume of the cube

OpenStudy (gabylovesyou):

volume of a cone = 1/3 volume of a cylinder

OpenStudy (gabylovesyou):

how is it different tho

OpenStudy (gabylovesyou):

hmmmm ? how r they different ?